A TV commercial claims that a car can speed up from rest to 60mph in 4.4s. If this is true, how many g's is the car's average acceleration?? There are 1,609m in a mile. Please explain thoroughly so I understand.
Answer #1

v = 60 x 1609 / 3600 = 26.82 m/s

v = u + at
26.82 = 0 + a*4.4
a = 6.09 m/s2

a =6.09 / 9.81 = 0.621 g